ANDY HAS 35 DIMES AND QUARTERS IN HIS PIGGY BANK WITH A TOTAL VALUE OF $6.95. HOW MANY OF THE COINS ARE DIMES AND HOW MANY ARE QUARTERS?
Answer by josmiceli(19441)

If he had 28 quarters, that's $7, so
he must have 27 quarters or less
27 quarters is $6.75 and that leaves $.20, or 2 dimes
27 + 2 = 29 (too small)
26 quarters = $6.50, leaving $.45 and then I need a nickel
25 quarters = $6.25 leaving $.70, or 7 dimes
25 + 7 = 32 (too small)
23 quarters = $5.75, leaving $1.20, or 12 dimes
23 + 12 = 35 Bingo
There are 23 quarters and 12 dimes
Notice that an even number of quarters won't work, since
that means he needs a nickel also